Find the slope if it exists:
x+5y=10

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

x+5y=10
5y=-x+10
y=(-x+10)/5
y=-1/5x+2
m=-1/5
y int = (0,2)

Answer by richwmiller(17219)

no need to do anything to the equation
in this form a=1 b=5
slope is -a/b=-1/5